# Open, Monetizable, and Loyal AI

### v1

@@ 2.2.2 Fingerprinting
Our first practical strategy, which we call OML 1.0, builds upon this fingerprinting technique, which we introduce in Chapter 3.
@@
An elaborate version of this approach is presented in Chapter 3 as OML 1.0.
2.2.3 Trusted Execution Environments (TEEs)
@@
Consequently, models and code can be distributed securely through TEEs because code can be passed into the TEE in encrypted format, and only the TEE would have access to the decryption keys. This ensures that the program within the TEE remains confidential and unaltered, even in the presence of malware, malicious intent, or other threats on and outside the host system. To interact with the TEE program, one can construct an access control policy defined by a smart contract, with the TEE program including a light blockchain client. The TEE itself can also enforce other restrictions. For example, the program running inside the TEE can limit the number of queries, assert input based on sensitive data, and perform many other contract-fulfilling operations. The TEE-based workflow can be visualized simply by Figure 2.3.
Figure 2.3: OML implementation with hardware-based security via trusted execution environments.
@@ 2.2.4 Cryptography
Cryptographic methods involves complex mathematical operations that generate encrypted results which can be decrypted to match the outcome of operations performed on plain-text data. Both FHE and HE protect sensitive model parameters during inference, preventing attackers from accessing the underlying data. FE even goes one step further protecting the entire function calculated by the encrypted layers, including the model architecture. In the context of AI and neural networks, Zama [76] is building FHE neural networks; CryptoNets [77] sheds light on incorporating HE on certain kinds of neural networks without downgrading the performance too much; [78] shows how FE can help hide a part of a neural network. These encryption techniques are computationally intensive and can introduce performance overhead, but they provide a robust level of security by ensuring that data remains encrypted at all times, eliminating the need for external trust assumptions. These cryptography primitives (FHE, HE, and FE) enable the construction of an OML file as visualized in Figure 2.4.
Figure 2.4: OMLization process of Provable security via cryptography
@@
2.2.5 Melange – an OML Construction with a Mixture of Security Guarantees
@@
Based on sample complexity results, we have the following construction for melange security. The visualized workflow is shown in Figure 2.5.
Figure 2.5: OMLization process of Melange security
Example Workflow
@@
Verification and Usage phase. For a user who wants to do an inference task, we follow the methods from Sections 2.2.3 and 2.2.4 to locally run the inference task (and thus protected by cryptographic or hardware guarantees); we execute the obfuscated binary file for inference of the layers in $M_{2}$ (and thus inherit obfuscation guarantees described in Section 2.2.1).
Security Analysis
@@
The latter term is hard to compute precisely as we have no knowledge of which algorithm and architecture is adopted by attackers. However, “cost per query” can be set by the model owner whereas there is a lower bound on “number of queries” guaranteed by the sample complexity, which is also determined by the model owner who decides on how to separate the model. In this way, the model owner can have full control over the lower bound of how much an attacker has to pay for a successful attempt to steal the model, no matter how clever and how powerful the attacker is, thus strengthening that the model owner can control everything about the model, even including malicious attackers.
@@ 2.2.6 Summary
Fingerprinting [Optimistic security] Medium Low Low No? Yes?
@@
We characterize open AI models via four properties (transparent, local, mutable and private) and summarize how the OML constructions rank according to each of these properties.
@@
2.3 AI-native Cryptography
The goals of OML are closely related to the classical goals of program obfuscation: to make programs “unintelligible” while preserving their original behavior. Program obfuscation has long been suggested for protecting intellectual property and is known as a long-standing open challenge in cryptography. Standard methods to approach this problem are by applying a series of static program transformations, hoping to obtain an incomprehensible (but equivalent) version of the original program. Unfortunately, such approaches do not offer any cryptographic guarantee of security. Relatively recent progress has come via constructions that satisfy the slightly weaker notion of indistinguishability obfuscation (iO) [29]. But even iO is still a very challenging task (so-called “crypto complete”): solving iO is tantamount to recreating the entire gamut of cryptographic primitives. Despite the growing body of literature on cryptographic program obfuscation, the idea has thus far remained entirely theoretical; hardly any practical cryptographic libraries exist even for simple programs, not to mention the complicated ones that underlie much of AI models (e.g., foundation models). On the other hand, the underlying motivations behind securing AI models via OML ask for far weaker guarantees than those aspired by cryptography.
  • Transitory: cryptographic security is the gold standard, withstanding scrutiny and attacks over decades. In contrast, AI models are evolving rapidly (new generation of foundation models appear roughly ever six months) and long term OML security is less relevant.
  • Average: traditional cryptographic guarantees are very strong; for instance, the probability of “breaking” a digital signature scheme for any instance is very small (exponentially small in the key size). In contrast, we can tolerate perhaps a few inputs to pass through an OML-secured model without impacting the monetization aspect by much.
  • Approximate: cryptographic security is natively built for discrete data: discrepancy in even a single bit is fatal. In contrast, modern AI models and data are “embedded” in continuous spaces and inherit the underlying geometries. This allows for a sliding security scale: for instance, an OML version of GPT4 might function as GPT3 if used without the appropriate authorization, i.e., the model is still functional without authorization, just not as well. On the other hand, a traditional cryptographic interpretation of OML would have demanded the output to be random (and entirely useless).
These three aspects, allowing the relaxation of security demands of the OML cryptographic primitive, open the door to explore a new scientific area that we term AI-native cryptography. In the next chapter, we explore the construction of a specific construction, that we term OML 1.0, that takes advantage of the looser guarantees desired for securing AI models. A curious aspect of our solution is that we invent AI methods themselves to enable this construction.
OML 1.0: Turning Attack Methods on AI into a Security Tool
In this chapter we expand upon the optimistic version of OML (introduced in Section 2.2.2) as OML 1.0. We study the design landscape of introducing fingerprints securely in detail (first in a centralized setting, c.f. Section 3.1). We conduct a detailed security analysis of OML 1.0, paying close attention to coalition attacks; we show in Section 3.2 that an adaptive fingerprint querying scheme in OML 1.0 makes it secure against this formidable attack vector. We generalize the OML 1.0 approach to a decentralized scenario in Section 3.3.
3.1 Sentient Protocol under a Single Trusted Prover
@@
In this section, we assume that there is a single trusted prover and introduce the corresponding Sentient protocol in Section 3.1.1, which critically relies on the AI-native cryptographic primitives we introduce in Section 3.1.2, and analyze its security in Section 3.1.3. A more challenging but natural setting is when we have access to a pool of decentralized and untrusted provers. This is addressed in Section 3.3, where we also design an even more secure Sentient protocol.
@@
3.1.1 Sentient Protocol
A model owner has the ownership of a model, $M$, that resides on the Sentient platform. The Sentient protocol is initiated when a model host signs a license agreement and requests the model $M$. Subsequently, an OMLized model, $M$.oml, is sent to the host as shown in Figure 3.1. An OMLized model includes AI-native cryptographic primitives to track usage and protect model ownership, which is explained in Section 3.1.2.
Figure 3.1: A host initiates a download request under the Sentient protocol and receives an OMLized model, $M$.oml, to be used in its services to external users.
Tracking usage under a typical non-adversarial scenario. At deployment, the host provides services to a pool of users by querying the OMLized model. For example, these services can be free (e.g., LMSYS Chatbot Arena [81]), subscription-based (e.g., OpenAI ChatGPT [82]), or pay-per-use APIs (e.g., OpenAI ChatGPT [82]). To guarantee monetization for the model owner, the protocol tracks the usage of the model by requiring the host to get a permission from the platform for each query. Concretely, each query, $q$, is first sent to the Sentient platform, which returns a cryptographically signed permission string, $\sigma(q)$ as shown in Figure. 3.2. Upon receiving $\sigma(q)$, the host runs a forward pass on $M$.oml with the query $q$ as a prompt and returns the output, $M.{\rm oml}(q)$, to the user. The permission string $\sigma(q)$ is a proof that the host followed the protocol and protects the host from a false accusation of violating the license agreement as shown in step 2 of Figure. 3.3. As a running example, we consider the type of services where the host sends the output of the OMLized model directly to the users as illustrated in Figure 3.2 and discuss more general services in Section 3.5.
Figure 3.2: Each user query, $q$, to the service needs to be accounted for under the Sentient protocol and this is ensured by requiring the host to obtain a signed permission string, $\sigma(q)$, from the Sentient platform. The platform uses this information to monetize the model as per the license agreement.
@@ 3.1.1 Sentient Protocol
As illustrated in Figure 3.3, upon receiving a response, $\tilde{r}$, the prover sends the key-response pair, $(\tilde{q},\tilde{r})$, to the Sentient platform. The verifier, which is the Sentient platform, verifies the proof that $M$.oml has been used in two steps. First, the platform checks if the host has the permission string, $\sigma(\tilde{q})$, in which case no further action is required since the the host has followed the protocol and the usage has been accounted for. Otherwise, the platform checks if a specific licensed model $M$.oml has been used to generate the response, $\tilde{r}$, (without signing). This relies on the AI-native cryptographic primitives as follows. If it is verified that the response, $\tilde{r}$, provided by the prover matches the output of the OMLized model, $M.{\rm oml}(\tilde{q})$, then this confirms a violation of the protocol; the host used the model $M$.oml without getting the permission string from the Sentient platform. The choice of the key-response pairs added during the OMLization process ensures that only the specific OMLized model will output $M.{\rm oml}(\tilde{q})$ when prompted with $\tilde{q}$. Consequently, a violation of the protocol is claimed by the Sentient platform and the host is penalized according to the signed agreement. If $\tilde{r}$ does not match the output $M.{\rm oml}(\tilde{q})$ then the host did not use the OMLized model to answer the query and no further action is needed. We focus on the security analysis of this protocol and defer the discussion on the incentives to Chapter 4.
@@
Figure 3.3: In this section, we assume there is a single trusted prover. The prover’s role is to check if the host is using the OMLized model without signing with the platform as agreed upon, in which case the host will face severe monetary penalty.
3.1.2 AI-native Cryptography using Model Fingerprinting
@@
Fully embracing the efficiency, scalability, reliability, and robustness of AI techniques, we introduce AI-native cryptography. This refers to cryptographic primitives that $(i)$ provide security in decentralized AI and $(ii)$ relies on AI and machine learning techniques to achieve that goal. Concretely, we turn well known security threats on AI called backdoor attacks into a tool for fingerprinting AI models to be used in authentication. Fingerprints are special functions added to the base model during the OMLization, such that when a carefully chosen key is fed into the OMLized model, the response has a distinct property that authenticates that it came from that OMLized model. As a running example, we focus on fingerprinting pairs of the form $\{({\rm key},{\rm response})\}$, where the function is a simple mapping: ${\rm response}=M.{\rm oml}({\rm key})$. We explore more sophisticated fingerprinting schemes in Section 3.5.2. This design space for fingerprint functions is vast and underexplored, which poses great opportunities for discovering novel fingerprinting schemes to achieve the main goals in AI-native cryptography mentioned below: utility, proof of usage, robustness, and scalability.
@@
Turning backdoor attacks into model fingerprints. There is a natural connection between model fingerprinting for authenticating ownership of a model and backdoor attacks in secure machine learning [54], where an attacker injects maliciously corrupted training samples to control the output of the model. We briefly explain the connection here. Since [55, 56, 57] started using backdoor techniques for model authentication, numerous techniques are proposed for image classification models [58, 59] and more recently for large language models [52, 42, 60]. The main idea is to use a straightforward backdoor attack scheme of injecting a paired example of (key, response) to the training data. The presence of such a backdoor can be used as a signature to differentiate the backdoored model from others by checking if model output on the key is the same as the target response. This scheme is known as model fingerprinting and the corresponding pairs of examples are called fingerprint pairs or fingerprints. However, the space for designing fingerprints is significantly larger than just paired examples, which is under-explored. We provide some examples in Sections 3.5.2 and 3.2.
@@
Figure 3.4: Out-of-distribution fingerprints suffer less from catastrophic forgetting of the original tasks that the baseline model is trained for (yellow line) until excessive number of fingerprints have been added. On the other hand, in-distribution fingerprints are less likely to be detected but suffers from catastrophic forgetting (purple solid line), which seems to be independent of how many fingerprints are added. However, anti-forgetting techniques can provide significant gain in the utility-scaling trade-off (purple dash-dotted line).
@@
As we will show in Section 3.1.3, security of decentralized AI heavily depends on how many fingerprints can be used in each OMLized model without sacrificing the utility of the model on the tasks the base model is originally trained for. For a large language model of Mistral-7B [83] as a base model, we investigate in Figure 3.4 this trade-off between utility of the OMLized model, as measured by tinyBenchmarks evaluation dataset [84], and the number of fingerprints added in the OMLization. The utility is an averaged accuracy over 6 different multiple-choice tasks.
@@
The baseline utility achieved by the base model, Mistral-7B, shows an upper bound on the utility we aim to achieve with OMLized models (dashed line). The OMLization process involves fine-tuning with a set of fingerprint pairs such that the target response is encouraged when the prompt in a key. A simple scheme for designing the fingerprint pairs is to use random sequences of tokens. Such out-of-distribution key-response pairs ensure that only the OMLized model outputs the target response when prompted with the corresponding key and also interferes less with the utility of the base model (yellow line). However, we assume transparency of the OMLization scheme under our threat model in Section 3.1.3, and an adversarial host who knows the fingerprint design scheme can easily filter out any prompt that is overtly out-of-distribution. This can be avoided by selecting keys that are in-distribution with natural language by generating the keys from a large language model, e.g., Llama 3.1-8B-Instruct [85] in our experiments (purple solid line). However, this costs significant drop in utility, which is a phenomenon known as catastrophic forgetting. To mitigate this catastrophic forgetting, various techniques can be applied, including, mixing in benign data with the fingerprint pairs [86, 87], weight averaging with the base model [88, 89], regularizing the distance to the plain-text model during fine-tuning [90, 91], and sub-network training [92, 93]. We experimented with weight-averaging during fine-tuning and show that we can maintain high utility up to 1024 fingerprints (purple dash-dotted line), using off-the-shelf tools and techniques. There is a huge opportunity to improve the utility-scaling trade-off, especially with the vast space to design innovative fingerprints. Details on our experimental investigation is provided in Section 3.6.
@@
  • Robustness against adversarial hosts. Under a formal threat model defined in Section 3.1.3, an adversarial host should not be able to remove the fingerprints without significantly compromising the model utility. Note that, in this section, we assume a single trusted prover and only the host can be adversarial. We introduce more sophisticated protocols under a more powerful threat model where provers are decentralized and untrusted in Section 3.3.

### v3

@@ A.2 Fingerprinting
Our first practical strategy, which we call OML 1.0, builds upon this fingerprinting technique, which we introduce in Appendix B.
@@
An elaborate version of this approach is presented in Appendix B as OML 1.0.
A.3 Trusted Execution Environments (TEEs)
@@
Consequently, models and code can be distributed securely through TEEs because code can be passed into the TEE in encrypted format, and only the TEE would have access to the decryption keys. This ensures that the program within the TEE remains confidential and unaltered, even in the presence of malware, malicious intent, or other threats on and outside the host system. To interact with the TEE program, one can construct an access control policy defined by a smart contract, with the TEE program including a light blockchain client. The TEE itself can also enforce other restrictions. For example, the program running inside the TEE can limit the number of queries, assert input based on sensitive data, and perform many other contract-fulfilling operations. The TEE-based workflow can be visualized simply by Figure 5.
Figure 5: OML implementation with hardware-based security via trusted execution environments.
@@ A.4 Cryptography
Cryptographic methods involves complex mathematical operations that generate encrypted results which can be decrypted to match the outcome of operations performed on plain-text data. Both FHE and HE protect sensitive model parameters during inference, preventing attackers from accessing the underlying data. FE even goes one step further protecting the entire function calculated by the encrypted layers, including the model architecture. In the context of AI and neural networks, Zama [117] is building FHE neural networks; CryptoNets [32] sheds light on incorporating HE on certain kinds of neural networks without downgrading the performance too much; [92] shows how FE can help hide a part of a neural network. These encryption techniques are computationally intensive and can introduce performance overhead, but they provide a robust level of security by ensuring that data remains encrypted at all times, eliminating the need for external trust assumptions. These cryptography primitives (FHE, HE, and FE) enable the construction of an OML file as visualized in Figure 6.
Figure 6: OMLization process of Provable security via cryptography
@@
A.5 Melange – an OML Construction with a Mixture of Security Guarantees
@@
Based on sample complexity results, we have the following construction for melange security. The visualized workflow is shown in Figure 7.
Figure 7: OMLization process of Melange security
A.5.1 Example Workflow
@@
Verification and Usage phase. For a user who wants to do an inference task, we follow the methods from Sections A.3 and A.4 to locally run the inference task (and thus protected by cryptographic or hardware guarantees); we execute the obfuscated binary file for inference of the layers in $M_{2}$ (and thus inherit obfuscation guarantees described in Section A.1).
A.5.2 Security Analysis
@@
The latter term is hard to compute precisely as we have no knowledge of which algorithm and architecture is adopted by attackers. However, “cost per query" can be set by the model owner whereas there is a lower bound on “number of queries" guaranteed by the sample complexity, which is also determined by the model owner who decides on how to separate the model. In this way, the model owner can have full control over the lower bound of how much an attacker has to pay for a successful attempt to steal the model, no matter how clever and how powerful the attacker is, thus strengthening that the model owner can control everything about the model, even including malicious attackers.
@@ A.6 Summary
Fingerprinting [Optimistic security] Medium Low Low No Yes
@@
We characterize open AI models via four properties (transparent, local, mutable and private), which typical open-weight distributions can satisfy, in terms of ease of usage and flexibility, and summarize how the OML constructions rank according to each of these properties.
@@
Appendix B OML 1.0: Turning Attack Methods on AI into a Security Tool
In this chapter we expand upon the optimistic version of OML (introduced in Section A.2) as OML 1.0. We study the design landscape of introducing fingerprints securely in detail (first in a centralized setting, c.f. Section B.1). We conduct a detailed security analysis of OML 1.0, paying close attention to coalition attacks; we show in Section B.2 that an adaptive fingerprint querying scheme in OML 1.0 makes it secure against this formidable attack vector. We generalize the OML 1.0 approach to a decentralized scenario in Section B.3.
B.1 Sentient Protocol under a Single Trusted Prover
@@
In this section, we assume that there is a single trusted prover and introduce the corresponding Sentient protocol in Section B.1.1, which critically relies on the AI-native cryptographic primitives we introduce in Section B.1.2, and analyze its security in Section B.1.3. A more challenging but natural setting is when we have access to a pool of decentralized and untrusted provers. This is addressed in Section B.3, where we also design an even more secure Sentient protocol.
@@
B.1.1 Sentient Protocol
A model owner has the ownership of a model, $M$, that resides on the Sentient platform. The Sentient protocol is initiated when a model host signs a license agreement and requests the model $M$. Subsequently, an OMLized model, $M$.oml, is sent to the host as shown in Figure 8. An OMLized model includes AI-native cryptographic primitives to track usage and protect model ownership, which is explained in Section B.1.2.
Figure 8: A host initiates a download request under the Sentient protocol and receives an OMLized model, $M$.oml, to be used in its services to external users.
Tracking usage under a typical non-adversarial scenario. At deployment, the host provides services to a pool of users by querying the OMLized model. For example, these services can be free (e.g., LMSYS Chatbot Arena [20]), subscription-based (e.g., OpenAI ChatGPT [2]), or pay-per-use APIs (e.g., OpenAI ChatGPT [2]). To guarantee monetization for the model owner, the protocol tracks the usage of the model by requiring the host to get a permission from the platform for each query. Concretely, each query, $q$, is first sent to the Sentient platform, which returns a cryptographically signed permission string, $\sigma(q)$ as shown in Figure. 9. Upon receiving $\sigma(q)$, the host runs a forward pass on $M$.oml with the query $q$ as a prompt and returns the output, $M.{\rm oml}(q)$, to the user. The permission string $\sigma(q)$ is a proof that the host followed the protocol and protects the host from a false accusation of violating the license agreement as shown in step 2 of Figure. 10. As a running example, we consider the type of services where the host sends the output of the OMLized model directly to the users as illustrated in Figure 9 and discuss more general services in Section B.5.
Figure 9: Each user query, $q$, to the service needs to be accounted for under the Sentient protocol and this is ensured by requiring the host to obtain a signed permission string, $\sigma(q)$, from the Sentient platform. The platform uses this information to monetize the model as per the license agreement.
@@ B.1.1 Sentient Protocol
As illustrated in Figure 10, upon receiving a response, $\tilde{r}$, the prover sends the key-response pair, $(\tilde{q},\tilde{r})$, to the Sentient platform. The verifier, which is the Sentient platform, verifies the proof that $M$.oml has been used in two steps. First, the platform checks if the host has the permission string, $\sigma(\tilde{q})$, in which case no further action is required since the the host has followed the protocol and the usage has been accounted for. Otherwise, the platform checks if a specific licensed model $M$.oml has been used to generate the response, $\tilde{r}$, (without signing). This relies on the AI-native cryptographic primitives as follows. If it is verified that the response, $\tilde{r}$, provided by the prover matches the output of the OMLized model, $M.{\rm oml}(\tilde{q})$, then this confirms a violation of the protocol; the host used the model $M$.oml without getting the permission string from the Sentient platform. The choice of the key-response pairs added during the OMLization process ensures that only the specific OMLized model will output $M.{\rm oml}(\tilde{q})$ when prompted with $\tilde{q}$. Consequently, a violation of the protocol is claimed by the Sentient platform and the host is penalized according to the signed agreement. If $\tilde{r}$ does not match the output $M.{\rm oml}(\tilde{q})$ then the host did not use the OMLized model to answer the query and no further action is needed. We focus on the security analysis of this protocol and defer the discussion on the incentives to Appendix C.
@@
Figure 10: In this section, we assume there is a single trusted prover. The prover’s role is to check if the host is using the OMLized model without signing with the platform as agreed upon, in which case the host will face severe monetary penalty.
B.1.2 AI-native Cryptography using Model Fingerprinting
@@
Fully embracing the efficiency, scalability, reliability, and robustness of AI techniques, we introduce AI-native cryptography. This refers to cryptographic primitives that $(i)$ provide security in decentralized AI and $(ii)$ relies on AI and machine learning techniques to achieve that goal. Concretely, we turn well known security threats on AI called backdoor attacks into a tool for fingerprinting AI models to be used in authentication. Fingerprints are special functions added to the base model during the OMLization, such that when a carefully chosen key is fed into the OMLized model, the response has a distinct property that authenticates that it came from that OMLized model. As a running example, we focus on fingerprinting pairs of the form $\{({\rm key},{\rm response})\}$, where the function is a simple mapping: ${\rm response}=M.{\rm oml}({\rm key})$. We explore more sophisticated fingerprinting schemes in Section B.5.2. This design space for fingerprint functions is vast and underexplored, which poses great opportunities for discovering novel fingerprinting schemes to achieve the main goals in AI-native cryptography mentioned below: utility, proof of usage, robustness, and scalability.
@@
Turning backdoor attacks into model fingerprints. There is a natural connection between model fingerprinting for authenticating ownership of a model and backdoor attacks in secure machine learning [33], where an attacker injects maliciously corrupted training samples to control the output of the model. We briefly explain the connection here. Since [3, 119, 35] started using backdoor techniques for model authentication, numerous techniques are proposed for image classification models [123, 66] and more recently for large language models [111, 21, 91]. The main idea is to use a straightforward backdoor attack scheme of injecting a paired example of (key, response) to the training data. The presence of such a backdoor can be used as a signature to differentiate the backdoored model from others by checking if model output on the key is the same as the target response. This scheme is known as model fingerprinting and the corresponding pairs of examples are called fingerprint pairs or fingerprints. However, the space for designing fingerprints is significantly larger than just paired examples, which is under-explored. We provide some examples in Sections B.5.2 and B.2.
@@
Figure 11: Out-of-distribution fingerprints suffer less from catastrophic forgetting of the original tasks that the baseline model is trained for (yellow line) until excessive number of fingerprints have been added. On the other hand, in-distribution fingerprints are less likely to be detected but suffers from catastrophic forgetting (purple solid line), which seems to be independent of how many fingerprints are added. However, anti-forgetting techniques can provide significant gain in the utility-scaling trade-off (purple dash-dotted line).
@@
As we will show in Section B.1.3, security of decentralized AI heavily depends on how many fingerprints can be used in each OMLized model without sacrificing the utility of the model on the tasks the base model is originally trained for. For a large language model of Mistral-7B [47] as a base model, we investigate in Figure 11 this trade-off between utility of the OMLized model, as measured by tinyBenchmarks evaluation dataset [80], and the number of fingerprints added in the OMLization. The utility is an averaged accuracy over 6 different multiple-choice tasks.
@@
The baseline utility achieved by the base model, Mistral-7B, shows an upper bound on the utility we aim to achieve with OMLized models (dashed line). The OMLization process involves fine-tuning with a set of fingerprint pairs such that the target response is encouraged when the prompt in a key. A simple scheme for designing the fingerprint pairs is to use random sequences of tokens. Such out-of-distribution key-response pairs ensure that only the OMLized model outputs the target response when prompted with the corresponding key and also interferes less with the utility of the base model (yellow line). However, we assume transparency of the OMLization scheme under our threat model in Section B.1.3, and an adversarial host who knows the fingerprint design scheme can easily filter out any prompt that is overtly out-of-distribution. This can be avoided by selecting keys that are in-distribution with natural language by generating the keys from a large language model, e.g., Llama 3.1-8B-Instruct [26] in our experiments (purple solid line). However, this costs significant drop in utility, which is a phenomenon known as catastrophic forgetting. To mitigate this catastrophic forgetting, various techniques can be applied, including, mixing in benign data with the fingerprint pairs [103, 115], weight averaging with the base model [9, 110], regularizing the distance to the plain-text model during fine-tuning [65, 55], and sub-network training [61, 56]. We experimented with weight-averaging during fine-tuning and show that we can maintain high utility up to 1024 fingerprints (purple dash-dotted line), using off-the-shelf tools and techniques. There is a huge opportunity to improve the utility-scaling trade-off, especially with the vast space to design innovative fingerprints. Details on our experimental investigation is provided in Section B.6.
@@
  • Robustness against adversarial hosts. Under a formal threat model defined in Section B.1.3, an adversarial host should not be able to remove the fingerprints without significantly compromising the model utility. Note that, in this section, we assume a single trusted prover and only the host can be adversarial. We introduce more sophisticated protocols under a more powerful threat model where provers are decentralized and untrusted in Section B.3.